# In-context Learning for Automated Driving Scenarios

### v1

@@ V-B RL Agent Selection and Implementation
DQN was chosen for our experimentation as it’s rather efficient and satisfied our research requirement. It shows suitability in handling discrete action spaces like AD problems and thus could be utilized to assess whether LLM could effectively guide RL in our scenario.
V-C Experiment Design and Performance Evaluation
@@
Mean reward, lane change percentage and speed up percentage are used as metrics to evaluate the performance of our agent. Mean reward represents the average total reward for actions taken in the testing scenario. Lane change percentage indicates the proportion of lane-changing actions out of the total actions, while speed-up percentage represents the proportion of acceleration actions out of the total actions. The agents are first trained in a customized environment. After training various agents, we conducted multiple experiments and analyses to assess their performance within a consistent and standardized highway testing environment.
Does LLM reward guide the training process of RL agents?
Using the sum of LLM reward and RL reward as a combined reward function for reinforcement learning and comparing it with the DQN baseline. We compare the performance by testing the average reward function obtained by running the trained agent in the same environment for 5 rounds. The results are as follows: The match rate increases through training aligning with the guidance of LLM. The match rate remains steady at around 50 per cent as shown in Fig. 5.
Figure 5: Match rate through training
Does prompt affect the performance or behaviour pattern of the LLM-RL agent?
Three agents were designed and implemented with different driving styles. The base driving agent, which utilizes the LLM’s general knowledge of driving without customized prompts, serves as the foundation for specialized adaptations. Specifically, the aggressive driving agent prioritizes efficiency, whereas the conservative driving agent emphasizes safety. The experiments were conducted in the same highway environment. Compared to the DQN baseline, both the conservative and base agents demonstrate relatively low lane-change and speed-up rates as shown in Fig. 6. This confirms that the behavioral patterns of these agents diverge from those of DQN agents, which frequently rush and make frequent lane changes to access higher rewards. Specifically, the aggressive agent displays a notably high speed-up rate compared to all agents, showcasing the influential effect of prompts.
(a) Collision Probabilities for different driving style agents
(b) Average speed for different driving style agents
Figure 6: Training Results
@@
Behaviour analysis in terms of mean reward, lane change and speed up are summarized in table I.
TABLE I: Experiment result: behaviour analysis
Index Mean Reward Lane Change Speed Up
@@
Detailed rewards are shown in Table IIbelow, confirming the previous conclusion. An aggressive agent denotes a rather high-speed reward compared to all other agents. Specifically, the highway-env reward can be broken down into four rewards in different aspects: collision reward, right lane reward, high-speed reward, and on road reward.
TABLE II: Experiment result: reward break down
Method Collision Lane Change High Speed On Road
Reward Reward Reward Reward
DQN baseline 0.02632 0.51316 0.29933 1.0
Aggressive 0.14706 0.41177 0.69912 1.0
Conservative 0.0 0.5 0.01435 1.0
Base 0.02255 0.58271 0.17517 1.0
The behaviour patterns of different agents are shown in the videos available at[34]. A conservative agent denotes a more stable behaviour and prioritises safety, which can be investigated by the 0 in collision reward. However, it also causes a lower speed referred to lower high speed reward. An aggressive agent shows a behavior pattern that prioritises efficiency. In this test case, the ego car takes frequent lane changes to maximize efficiency, to a certain extent, disregarding safety.
VI Conclusion

### v2

@@ V-B RL Agent Selection and Implementation
DQN was chosen for our experimentation as it is rather efficient and satisfied our research requirement. It shows suitability in handling discrete action spaces like AD problems and thus could be utilized as a baseline to assess whether LLM could effectively guide RL in our scenario.
V-C Evaluation criteria definition
Collision Score
The collision score penalizes the agent for causing collisions. It is calculated based on whether a collision occurs during the agent’s driving, with a higher penalty for more severe collisions.
$S_{\text{collision}}(s,a)=-\text{Collision}(s,a)$ (9)
where $\text{Collision}(s,a)$ is a binary indicator (1 for collision, 0 for no collision).
Lane Change Score
The lane change score encourages or discourages lane changes based on driving behavior. Excessive lane changes may lead to a lower score, while appropriate lane changes can achieve a higher score.
$S_{\text{lane\_change}}(s,a)=\text{LaneChange}(s,a)$ (10)
where $\text{LaneChange}(s,a)$ is the number of lane changes performed in the state $s$ with action $a$.
High Speed Score
The high-speed reward encourages the agent to maintain a high speed while staying within safe limits. The reward increases as the agent approaches the target speed but decreases if the speed exceeds the defined safety threshold.
$S_{\text{high\_speed}}(s,a)=\left(\frac{v(s,a)-v_{\text{min}}}{v_{\text{desired}}-v_{\text{min}}}\right)$ (11)
where $v(s,a)$ is the agent’s speed in state $s$ with action $a$, and $v_{\text{desired}}$ is the optimal speed.
On Road Score
The on-road score keeps the agent within designated lanes, awarding a consistent score as long as it stays within the correct boundaries.
$S_{\text{on\_road}}(s,a)=\text{OnRoad}(s,a)$ (12)
where $\text{OnRoad}(s,a)$ is a binary indicator (1 if on road, 0 otherwise).
VI Performance Evaluation
@@
To evaluate the performance of the agents, we used three key metrics: mean reward, lane change percentage, and speed-up percentage. Mean reward represents the average total reward for actions taken in the testing scenario. Lane change percentage indicates the proportion of lane-changing actions out of the total actions, while speed-up percentage represents the proportion of acceleration actions out of the total actions. The agents were trained in a customized environment before conducting experiments in a consistent and standardized highway testing environment.
VI-A Does LLM reward guide the training process of RL agents?
We combined LLM and RL rewards to train agents and compared their performance against the DQN baseline. The average reward was calculated over five testing rounds in the same environment. As shown in Fig. 5, the match rate consistently improved during training, reflecting the influence of LLM guidance. The match rate stabilized at approximately 50%, demonstrating the agent’s alignment with LLM-provided instructions.
Figure 5: Match rate through training.
(a) Collision probabilities for different driving styles.
(b) Average speed for different driving styles.
Figure 6: Training results for driving style agents.
VI-B Does prompt affect the performance or behavior pattern of the LLM-RL agent?
Three agents with different driving styles were implemented:
  • Base Agent: Uses the LLM’s general knowledge of driving without specific customization.
  • Aggressive Agent: Prioritizes efficiency and speed.
  • Conservative Agent: Prioritizes safety and stability.
The experiments were conducted in the same highway environment. As shown in Fig. 6, compared to the DQN baseline, the conservative agent demonstrated the lowest lane-change and speed-up rates, while the aggressive agent displayed the highest speed-up rate, reflecting the influence of prompts in shaping behavior.
VI-C Behavior Analysis
@@
Behavior analysis, including mean reward, lane changes, and acceleration rates, is summarized in Table I.
TABLE I: Experiment result: behavior analysis including PPO baseline.
Index Mean Score Lane Change Score Speed Up Score
@@
BC-SAC [33] 0.83410 0.01750 0.75530
LLM-RL 0.84532 0.01045 0.81233
The aggressive agent achieves the highest speed-up score, while the conservative agent demonstrates minimal lane-change and speed-up actions. The LLM-RL agent achieves the highest overall mean score, showcasing a balance between safety and efficiency.
VI-D How does the reward ablation affect the performance?
To investigate the impact of different reward components on the agent’s behavior, we conducted a reward ablation study. The results are summarized in Table II.
TABLE II: Reward ablation study: reward breakdown.
Configuration Collision Score Lane Change Score High Speed Score
Safety Only -0.05 0.23 0.18
Efficiency Only -0.20 0.48 0.72
LLM Only -0.10 0.31 0.33
Safety + Efficiency -0.12 0.35 0.55
Safety + LLM -0.08 0.28 0.42
Efficiency + LLM -0.15 0.41 0.63
All (Full Reward) -0.12 0.53 0.75
VI-E Analysis of Results
1. Collision Score: The Safety Only configuration achieves the least negative Collision Score ($-0.05$), indicating the lowest collision risk. The Efficiency Only configuration has the most negative Collision Score ($-0.20$), reflecting higher collision risk due to its aggressive behavior. The All (Full Reward) configuration balances safety and efficiency, achieving a Collision Score of $-0.12$, which is better than Efficiency Only and closer to Safety Only. 2. Lane Change Score: The Efficiency Only and All (Full Reward) configurations show the highest Lane Change Scores ($0.48$ and $0.53$, respectively), indicating frequent lane-changing behavior. The Safety Only configuration has the lowest Lane Change Score ($0.23$), reflecting more conservative driving. 3. High Speed Score: The Efficiency Only configuration achieves the highest High Speed Score ($0.72$), emphasizing speed as the primary goal. The Safety Only configuration has the lowest High Speed Score ($0.18$). The All (Full Reward) configuration performs close to Efficiency Only ($0.75$) while maintaining a balanced Collision Score. The reward ablation study demonstrates that:
  • Single-reward configurations (e.g., Safety Only or Efficiency Only) lead to extreme behaviors, such as overly conservative or aggressive driving.
  • The All (Full Reward) configuration achieves the best balance, delivering high efficiency with acceptable safety performance.
  • Incorporating LLM rewards significantly enhances decision-making, aligning agent behavior with human-like patterns.
This analysis highlights the importance of designing comprehensive reward structures to optimize performance in dynamic environments while maintaining safety.
VII Conclusion